ADNAN MUHAMMED

Grade 12,

A special kind of light bulb emits monochromatic light of wavelength 700nm. Electrical energy supply to it at the rate of 60W and the bulb is 50% efficient at converting that energy to light energy. How many photons are emitted by the bulb during it`s life time of 1 day? (Give me solution using formula )

A special kind of light bulb emits monochromatic light of wavelength 700nm. Electrical energy supply to it at the rate of 60W and the bulb is 50% efficient at converting that energy to light energy. How many photons are emitted by the bulb during it`s life time of 1 day? (Give me solution using formula )

Grade:11

1 Answers

Arun
6 years ago
Dear Preeti
 
As we know
n *h* \nu = P
n = P/h\nu = P\lambda/ hc
On putting the values
n = 2.1 * 10^20 photons/sec
Since efficiency = 50%
Hence
50% of n = 1.05 * 10^20 photons/sec
Lifetime is 1 day
Hence nos of photons = 1.05 *10^20 * 24*60*60
 = 9 * 10^24 photons per day

Regards
Arun (askIITians forum expert)